The 18-month, and 24-month risk-free zero rates are 4.75%, and 5% with semiannual compounding. What is the value of an FRA where the holder pays LIBOR and receives 7% (semiannually compounded) for a six-month period beginning in 18 months? The principak is $10 million
Answer #1

0v18 = 4.75%

0v24 = 5.00%

18v24 = (1.0500)4/(1.0475)3 = 1.0575 or 5.75%

Payment = $10,000,000 * 5.75% * 6/12 = $287,500

Receipt = $10,000,000 * 7.00% * 6/12 = $350,000

The value of the FRA will be the difference between the receipt and payment

Value of FRA = $350,000 - $287,500 = $62,500